### v3

@@ 4 Upper Bounds
It is easy to see that the drawing is upward since it is consistent with the linear extension. Consider an arbitrary edge $(u,v)$ with $x(u)<x(v)$. The length of any edge in x-direction is bounded by $\operatorname{bw}(G)$; let $x(v)-x(u)=\ell\leq\operatorname{bw}(G)$. Edge $(u,v)$ may be crossed (a) by edges that have at least one incident vertex $w$ with $x(u)<x(w)<x(v)$, or (b) by edges $(w,w^{\prime})$ with $x(w)<x(u)\wedge x(v)<x(w^{\prime})$. Thus, there are at most $\ell-1$ and $\operatorname{bw}(G)-\ell-1$ possible choices for $w$ in the two scenarios, respectively. Since each vertex is incident to at most $\Delta$ edges, $(u,v)$ can be crossed at most $\Delta(\operatorname{bw}(G)-2)$ times.
For some graphs, a sublinear bound on the bandwidth is known, see [12, 22, 43]. This gives upper bounds on the upward local crossing number of many graph classes. We list only a few:
Corollary 4.3.
@@
  • Planar graphs of maximum degree $\Delta$ have bandwidth ${\mathcal{O}}(\frac{n}{\log_{\Delta}n})$ [12], hence their upward local crossing number is in $\mathcal{O}(\frac{n\cdot\Delta}{\log_{\Delta}n})$.
We complement the negative result in Proposition 3.1 by showing that every directed acyclic outerpath allows an upward 2-planar drawing. We start with a technical lemma on fans.
@@
Figure 6: Constructing upward $2$-planar drawings of fans according to Lemma 4.4.
Lemma 4.4.
@@
Proof 4.5.
We place $c$ at $(n-1,0)$; then we place $v_{1},v_{2},\dots,v_{n-1}$ above or below $c$ depending on the direction of the edges that connect them to $c$; see Figure 6 for an example.
@@
We now consider the case that $j=n$ or $(v_{j-1},v_{j-2})\in E(G)$; see, for example, path $P_{5}$ in Figure 6. In this case, we place $v_{h+1},\dots,v_{j-1}$ using x-coordinates $h+1,\dots,j-1$, going up and down as needed but remaining below the x-axis. The edges are drawn such that all vertices of $P^{\prime}$ remain on the outer face of the drawing. I.e., if we use straight-line edges, then, for $i\in[n-2]$, the slope of $v_{i}v_{i+1}$ must be less than the slope of $v_{i}c$. Since we go towards $c$, we can draw $P^{\prime}$ and the edges that connect $v_{1},v_{2},\dots,v_{n-1}$ to $c$ without any crossings.
If $j\neq n$ and $(v_{j-2},v_{j-1})\in E(G)$, then let $i\in\{h,\dots,j-1\}$ be the smallest index such that the subpath $\langle v_{i},v_{i+1},\dots,v_{j-1}\rangle$ is directed. See, for example, that last part of $P_{1}$ in Figure 6. In that case, we place $v_{h+1},\dots,v_{i-1}$ at x-coordinates $h+1,\dots,i-1$, going slightly up and down as in the case described above. Let $y_{\min}$ be the smallest among the y-coordinates of all points placed so far.
Then we place $v_{i},v_{i+1},\dots,v_{j-1}$ in reverse order, i.e., at x-coordinates $j-1,j-2,\dots,i$. For the y-coordinates, we choose $y(v_{i})=y_{\min}-\gamma$ and $y(v_{j-1})=y_{\min}-\varepsilon$ for some (large) $\gamma>0$ and (small) $\varepsilon>0$ with the following properties: if $i>h$ then $v_{j-1}$ lies inside the triangle $\triangle v_{i-1}v_{i}c$ (pale yellow in Figure 6) and if $i=h$ then $v_{j-1}$ lies inside the triangle $\triangle ov_{i}c$, where $o=(0,0)$. Draw $v_{i+1},\dots,v_{j-2}$ on the segment $\overline{v_{i}v_{j-1}}$. This fulfills the invariant: if $i>h$ then the vertex $v_{j-1}$ can reach the outer face via the edge $(v_{i},v_{i-1})$ which was not crossed so far. If $i=h$ then $v_{j-1}$ is on the outer face if $P^{\prime}=P_{1}$, otherwise $v_{j-1}$ can reach the current outer face by crossing the edge $(v_{h},v_{h-1})$. Observe that, by our invariant, $(v_{h},v_{h-1})$ might have crossed one edge in order to reach the outer face. While drawing $P^{\prime}$, we do not cross $(v_{h},v_{h-1})$ again. So the potential crossing with $(v_{j-1},v_{j})$ is at most the second crossing of $(v_{h},v_{h-1})$, and $(v_{h},v_{h-1})$ will not be crossed later.
Observe that when we draw the next maximal subpath, we place $v_{j}$ at $(j,y_{\max}+1)$, i.e., in particular in the outer face of the current drawing. The edge from $v_{j-1}$ to $v_{j}$ must be directed towards $v_{j}$ since the orientation is acylic. Thus, we can draw the edge between $v_{j-1}$ and $v_{j}$ upward with at most one crossing, causing at most a second crossing on $(v_{h},v_{h-1})$.
@@
Figure 7: Example in- and output of our drawing algorithm (edge crossings due to Lemma 4.4 are highlighted in yellow; other edge crossings are highlighted in orange).
Now we describe our construction for general outerpaths; see Figure 7.
Theorem 4.6.
@@
Let $G^{\prime}$ be such a graph; see Figure 7. Let $c_{1},c_{2},\dots,c_{k}$ be the vertices of degree at least 4 in $G^{\prime}$ (marked red in Figure 7). These vertices form a path (light red in Figure 7); let them be numbered along this path, which we call the backbone of $G^{\prime}$. We assign every vertex $v$ that does not lie on the backbone to a neighboring backbone vertex; if $v$ is incident to an inner edge, we assign $v$ to the other endpoint of that edge. Otherwise $v$ has degree 2 and is incident to a unique backbone vertex via an outer edge, and we assign $v$ to this backbone vertex. For $i\in[k]$, backbone vertex $c_{i}$ induces, together with the vertices assigned to it, a fan $F_{i}$.
@@
We draw the backbone in an x-monotone fashion. We start by drawing $F_{1}$ with the algorithm for drawing a fan as detailed in the proof of Lemma 4.4; see the leftmost gray box in Figure 7. Then, for $i\in\{2,\dots,k\}$, we set $x(c_{i})$ to $x(c_{i-1})$ plus the number of inner edges incident to $c_{i}$ and we set $y(c_{i})$ depending (i) on the y-coordinates of the two neighbors of $c_{i}$ that have already been drawn ($c_{i-1}$ and the common neighbor $u_{i-1}$ of $c_{i-1}$ and $c_{i}$ in $F_{i-1}$) and (ii) on the directions of the edges that connect these vertices to $c_{i}$; see, for example, the placement of $c_{5}$ in Figure 7. Then we draw $F_{i}$ with respect to the position of $c_{i}$, again using the algorithm from the proof of Lemma 4.4 with the following modifications. In general, vertices in $F_{i}$ that are adjacent to $c_{i}$ via an edge directed towards $c_{i}$ (resp. from $c_{i}$) are placed below (resp. above) all vertices in the drawings of $F_{1},\dots,F_{i}$; see the dark gray boxes below (resp. above) $c_{2},\dots,c_{5}$ in Figure 7. If an edge of $F_{i}$ connects two neighbors of $c_{i}$ one of which lies above $c_{i}$ and one of which lies below $c_{i}$, then we route this edge to the left of all drawings of $F_{1},\dots,F_{i-1}$.
@@
An exception to this rule occurs if $c_{i}$ and the common neighbor $w_{i}$ of $c_{i-1}$ and $c_{i}$ in $F_{i}$ must be both above or both below $c_{i-1}$ due to the directions of the corresponding edges. Let $u_{i-1}$ be the common neighbor of $c_{i-1}$ and $c_{i}$ in $F_{i-1}$. We assume, without loss of generality, that $c_{i}$ is above $c_{i-1}$. Let $P_{1}$ and $P_{2}$ be the first and second maximal subpath from Lemma 4.4 applied to $F_{i}$, and let $e_{i}$ be the edge connecting $P_{1}$ and $P_{2}$. We distinguish two subcases.
@@
If $P_{1}$ is a directed path leaving $w_{i}$, then we draw $P_{1}$ above the edge $c_{i-1}c_{i}$ and we draw the edge $e_{i}$ straight, without going around all drawings of $F_{1},\dots,F_{i-1}$. In this case $e_{i}$ is directed from $P_{1}$ to $P_{2}$. Hence, $e_{i}$ crosses the edge $u_{i-1}c_{i}$ if $u_{i-1}c_{i}$ is directed from $c_{i}$ to $u_{i-1}$; see the situation for $c_{2}$ in Figure 7. Note that by Property 4 of Lemma 4.4, $e_{i}$ may receive at most a second crossing when we draw the remainder of $F_{i}$ in the usual way.
@@
Otherwise, that is, if $P_{1}$ contains an edge directed towards the firstendpoint $w_{i}$ of $P_{1}$, let $f$ be the first such edge. We then place the part of $P_{1}$ up to the first endpoint of $f$ below the edge $c_{i-1}c_{i}$; see $w_{5}$ and $f$ in Figure 7. If the edge $u_{i-1}c_{i}$ is directed towards $c_{i}$, we draw it between $w_{i}$ and the edge $c_{i-1}c_{i}$. Then it crosses the edge $c_{i-1}w_{i}$ but no other edge. We place the second endpoint of $f$ below all vertices in $V(F_{1})\cup\dots\cup V(F_{i-1})$ and continue with the remainder of $F_{i}$ as usual.
@@
In any case, if $1<i<k$, then the last vertex $u_{i-1}$ of $F_{i-1}$ is connected to $c_{i}$ and $c_{i-1}$ is connected to the first vertex $w_{i}$ in $F_{i}$. These two edges may cross each other; see the crossings highlighted in orange in Figure 7. If the edge $c_{i-1}w_{i}$ goes, say, up but the following outer edges go down until a vertex $v_{k}$ below $c_{i}$ is reached, then the edge $c_{i-1}w_{i}$ may be crossed a second time by the edge $v_{k-1}v_{k}$; see the crossing labeled $x$ on the edge $c_{3}w_{4}$ in Figure 7. Observe that in this case the path $P_{1}$ is the directed path from $w_{i}$ to $v_{k-1}$. Thus, due to Property 4 of Lemma 4.4,edge $v_{k-1}v_{k}$ had been crossed at most once within its fan. Also $c_{i-1}w_{i}$ cannot have a third crossing. Thus, all edges are crossed at most twice.
@@
One can argue that every maximal pathwidth-2 graph can be generated from a maximal outerpath by connecting some pairs of adjacent vertices using an arbitrary number of (new) paths of length 2. In spite of the simplicity of this operation, we cannot hope to generalize the above result to pathwidth-2 graphs; see the linear lower bound on the upward local crossing number for such graphs stated in Theorem 3.5.
@@ 5 Testing Upward 1-Planarity
We start with a definition. Let $G_{1}$ and $G_{2}$ be any two st–digraphs. Let $s_{i}$ be the source and $t_{i}$ be the sink of $G_{i}$, with $i=1,2$. Let $G$ be a digraph that contains both $G_{1}$ and $G_{2}$ as induced subgraphs. Let $\Gamma$ be a drawing of $G$ and let $\Gamma_{1,2}$ be the drawing obtained by restricting $\Gamma$ to the vertices and edges of $G_{1}\cup G_{2}$. We say that $G_{1}$ and $G_{2}$ fully cross in $\Gamma$ if in $\Gamma_{1,2}$ every $s_{1}t_{1}$-path (i.e., a path directed from $s_{1}$ to $t_{1}$) crosses every $s_{2}t_{2}$-path. See Figures 8(a) and 8(b) for examples of st-digraph $G_{1}$ and $G_{2}$ that do not fully cross or fully cross in a drawing of $\Gamma_{1,2}$, respectively.
@@
We now define a few gadgets; all of them are planar st-graphs. For positive integers $b$ and $q$, let a $(b,q)$-parallel be the parallel composition of $b$ oriented paths each consisting of $q$ edges; see Figure 9(a). For a positive integer $p$, let a $(p)$-gate be the parallel composition of an oriented edge and a $(p-1,2)$-parallel; see Figure 9(b). For positive integers $h$, $q$, and $a$, let an $(h,q,a)$-chain consist of a series of $h$ $(q)$-gates, followed by exactly one $(a)$-gate, followed again by $h$ $(q)$-gates; see Figure 9(c).
@@
Conversely, assume that the instance $I$ of 3-Partition admits a solution. In order to prove that $G$ admits an upward 1-planar drawing, by hypothesis (ii) it suffices to construct an upward 1-planar drawing $\Gamma_{AB}$ of $G_{A}$ and $G_{B}$ where $G_{A}$ and $G_{B}$ fully cross. We enumerate the paths of $G_{B}$ as $\pi_{1},\pi_{2},\dots,\pi_{b}$. We also enumerate the $2(b-1)$ many $(W+1)$-gates and the $(a_{i})$-gate of a $(b-1,W+1,a_{i})$-chain as $g_{i,1},g_{i,2},\cdots,g_{i,2(b-1)}$, where $g_{i,1}$ contains the source $s_{A}$ of $G_{A}$ and $g_{i,2(b-1)}$ contains the sink $t_{A}$ of $G_{A}$ (see also Figure 10).Consider a path $\pi_{j}$ with $1\leq j\leq b$, and let $\{a_{\chi},a_{\lambda},a_{\mu}\}$ be the $j$-th bin of the solution of $I$. Let $\nu$ be an index such that $1\leq\nu\leq b$.
  • If $\nu$ is one of $\chi,\lambda,\mu$, then $\pi_{j}$ crosses the $(a_{\nu})$-gate of the $(b-1,W+1,a_{\nu})$-chain (see, for example, the red path crossing $g_{1,3}$ in Figure 10).
  • If $\nu$ is not one of $\chi,\lambda,\mu$ and the path $\pi_{h}$ that crosses the $(a_{\nu})$-gate of the $(b-1,W+1,a_{\nu})$-chain is such that $h>j$, then $\pi_{j}$ crosses the gate $g_{\nu,j}$ of the $(W+1,b-1,a_{\nu})$-chain (see, for example, the blue path crossing $g_{2,2}$ in Figure 10).
  • Otherwise ($h<j$), path $\pi_{j}$ crosses the gate $g_{\nu,b-1+j}$ of the $(b-1,W+1,a_{\nu})$-chain (see, for example, the red path crossing $g_{5,5}$ in Figure 10).
By the procedure above, if $\pi_{j}$ crosses a gate $g(\nu,q)$, with $1\leq\nu\leq k$ and $1\leq q\leq 2(b-1)$, there is no other $\pi_{h}$, with $h\neq j$, such that $\pi_{h}$ crosses $g(\nu,q)$. Also, the number of edges that $\pi_{j}$ crosses is $a_{\chi}+a_{\lambda}+a_{\mu}+(k-3)(W+1)=W+(k-3)(W+1)$, which is the number of edges of $\pi_{j}$. Hence, it is possible to draw $\pi_{j}$ in such a way that each of its edges is crossed exactly once and no edge of $G_{A}$ is crossed more than once. This concludes the proof.
@@
Theorem 5.3.
@@
Proof 5.4.
It is immediate to observe that upward 1-planarity testing is in the NP class of complexity (one can guess an upward embedding and test it in polynomial time). We show that it is NP-hard. For each case in the statement it suffices to exhibit a digraph $G$ that contains $G_{A}$ and $G_{B}$ as sub-graphs and that satisfies the conditions of Theorem 5.1.

### v4

@@ 4 Upper Bounds
It is easy to see that the drawing is upward since it is consistent with the linear extension. Consider an arbitrary edge $(u,v)$ and assume, without loss of generality, that $x(u)<x(v)$. Let $\ell=x(v)-x(u)$. The length of any edge in x-direction is bounded by $\operatorname{bw}(G)$; hence $\ell\leq\operatorname{bw}(G)$. Edge $(u,v)$ may be crossed (a) by edges that have at least one incident vertex $w$ with $x(u)<x(w)<x(v)$, or (b) by edges between two vertices $w$ and $w^{\prime}$ with $x(w)<x(u)$ and $x(v)<x(w^{\prime})$. Thus, there are at most $\ell-1$ and $\operatorname{bw}(G)-\ell-1$ possible choices for $w$ in the two scenarios, respectively. Since each vertex is incident to at most $\Delta$ edges, $(u,v)$ can be crossed at most $\Delta(\operatorname{bw}(G)-2)$ times. ∎
For some graph classes, sublinear bounds on the bandwidth are known (see, e.g., Böttcher, Pruessmann, Taraz, and Würfl, 2010; Fomin and Golovach, 2003; Wood, 2006). This gives upper bounds on the upward local crossing number of many graph classes. We list only a few:
Corollary 2.
@@
  • Planar graphs of maximum degree $\Delta$ have bandwidth ${\mathcal{O}}({n}/{\log_{\Delta}n})$ (Böttcher et al., 2010), hence their upward local crossing number is in $\mathcal{O}({n\cdot\Delta}/{\log_{\Delta}n})$. Observe that ${n\cdot\Delta}/{\log_{\Delta}n}\in o(n)$ if $\Delta\in o({\log n}/{\log\log n})$.
We now complement the negative result in Proposition 1 by showing that every directed acyclic outerpath allows an upward 2-planar drawing. We start with a technical lemma on fans.
@@
Figure 6: Constructing upward $2$-planar drawings of fans according to Lemma 1.
Lemma 1.
@@
Proof.
We place $c$ at $(n-1,0)$; then we place $v_{1},v_{2},\dots,v_{n-1}$ above or below $c$ depending on the direction of the edges that connect them to $c$; see Fig. 6 for an example.
@@
We now consider the case that $j=n$ or $(v_{j-1},v_{j-2})\in E(G)$; see, for example, path $P_{5}$ in Fig. 6. In this case, we place $v_{h+1},\dots,v_{j-1}$ using x-coordinates $h+1,\dots,j-1$, going up and down as needed but remaining below the x-axis. The edges are drawn such that all vertices of $P^{\prime}$ remain on the outer face of the drawing. I.e., if we use straight-line edges, then, for $i\in[n-2]$, the slope of $v_{i}v_{i+1}$ must be less than the slope of $v_{i}c$. Since we go towards $c$, we can draw $P^{\prime}$ and the edges that connect $v_{1},v_{2},\dots,v_{n-1}$ to $c$ without any crossings.
If $j\neq n$ and $(v_{j-2},v_{j-1})\in E(G)$, then let $i\in\{h,\dots,j-1\}$ be the smallest index such that the subpath $\langle v_{i},v_{i+1},\dots,v_{j-1}\rangle$ is directed. See, for example, that last part of $P_{1}$ in Fig. 6. In that case, we place $v_{h+1},\dots,v_{i-1}$ at x-coordinates $h+1,\dots,i-1$, going slightly up and down as in the case described above. Let $y_{\min}$ be the smallest among the y-coordinates of all points placed so far.
Then we place $v_{i},v_{i+1},\dots,v_{j-1}$ in reverse order, i.e., at x-coordinates $j-1,j-2,\dots,i$. For the y-coordinates, we choose $y(v_{i})=y_{\min}-\gamma$ and $y(v_{j-1})=y_{\min}-\varepsilon$ for some (large) $\gamma>0$ and (small) $\varepsilon>0$ with the following properties: if $i>h$ then $v_{j-1}$ lies inside the triangle $\triangle v_{i-1}v_{i}c$ (pale yellow in Fig. 6) and if $i=h$ then $v_{j-1}$ lies inside the triangle $\triangle ov_{i}c$, where $o=(0,0)$. Draw $v_{i+1},\dots,v_{j-2}$ on the segment $\overline{v_{i}v_{j-1}}$. This fulfills the invariant: if $i>h$ then the vertex $v_{j-1}$ can reach the outer face via the edge $(v_{i},v_{i-1})$ which was not crossed so far. If $i=h$ then $v_{j-1}$ is on the outer face if $P^{\prime}=P_{1}$, otherwise $v_{j-1}$ can reach the current outer face by crossing the edge $(v_{h},v_{h-1})$. Observe that, by our invariant, $(v_{h},v_{h-1})$ might have crossed one edge in order to reach the outer face. While drawing $P^{\prime}$, we do not cross $(v_{h},v_{h-1})$ again. So the potential crossing with $(v_{j-1},v_{j})$ is at most the second crossing of $(v_{h},v_{h-1})$, and $(v_{h},v_{h-1})$ will not be crossed later.
Moreover, when we draw the next maximal subpath, we place $v_{j}$ at $(j,y_{\max}+1)$, i.e., in particular in the outer face of the current drawing. The edge from $v_{j-1}$ to $v_{j}$ must be directed towards $v_{j}$ since the orientation is acylic. Thus, we can draw the edge between $v_{j-1}$ and $v_{j}$ upward with at most one crossing, causing at most a second crossing on $(v_{h},v_{h-1})$. ∎
@@
Figure 7: Example in- and output of our drawing algorithm (edge crossings due to Lemma 1 are highlighted in yellow; other edge crossings are highlighted in orange).
Now we describe our construction for general outerpaths; see Fig. 7.
Theorem 4.
@@
Let $G^{\prime}$ be such a graph; see Fig. 7. Let $c_{1},c_{2},\dots,c_{k}$ be the vertices of degree at least 4 in $G^{\prime}$ (marked red in Fig. 7). These vertices form a path (light red in Fig. 7); let them be numbered along this path, which we call the backbone of $G^{\prime}$. We assign every vertex $v$ that does not lie on the backbone to a neighboring backbone vertex; if $v$ is incident to an inner edge, we assign $v$ to the other endpoint of that edge. Otherwise $v$ has degree 2 and is incident to a unique backbone vertex via an outer edge, and we assign $v$ to this backbone vertex. For $i\in[k]$, backbone vertex $c_{i}$ induces, together with the vertices assigned to it, a fan $F_{i}$.
@@
We draw the backbone in an x-monotone fashion. We start by drawing $F_{1}$ with the algorithm for drawing a fan as detailed in the proof of Lemma 1; see the leftmost gray box in Fig. 7. Then, for $i\in\{2,\dots,k\}$, we set $x(c_{i})$ to $x(c_{i-1})$ plus the number of inner edges incident to $c_{i}$ and we set $y(c_{i})$ depending (i) on the y-coordinates of the two neighbors of $c_{i}$ that have already been drawn ($c_{i-1}$ and the common neighbor $u_{i-1}$ of $c_{i-1}$ and $c_{i}$ in $F_{i-1}$) and (ii) on the directions of the edges that connect these vertices to $c_{i}$; see, for example, the placement of $c_{5}$ in Fig. 7. Then we draw $F_{i}$ with respect to the position of $c_{i}$, again using the algorithm from the proof of Lemma 1 with the following modifications. In general, vertices in $F_{i}$ that are adjacent to $c_{i}$ via an edge directed towards $c_{i}$ (resp. from $c_{i}$) are placed below (resp. above) all vertices in the drawings of $F_{1},\dots,F_{i}$; see the dark gray boxes below (resp. above) $c_{2},\dots,c_{5}$ in Fig. 7. If an edge of $F_{i}$ connects two neighbors of $c_{i}$ one of which lies above $c_{i}$ and one of which lies below $c_{i}$, then we route this edge to the left of all drawings of $F_{1},\dots,F_{i-1}$.
@@
An exception to this rule occurs if $c_{i}$ and the common neighbor $w_{i}$ of $c_{i-1}$ and $c_{i}$ in $F_{i}$ must be both above or both below $c_{i-1}$ due to the directions of the corresponding edges. Let $u_{i-1}$ be the common neighbor of $c_{i-1}$ and $c_{i}$ in $F_{i-1}$. We assume, without loss of generality, that $c_{i}$ is above $c_{i-1}$. Let $P_{1}$ and $P_{2}$ be the first and second maximal subpath from Lemma 1 applied to $F_{i}$, and let $e_{i}$ be the edge connecting $P_{1}$ and $P_{2}$. We distinguish two subcases.
@@
If $P_{1}$ is a directed path leaving $w_{i}$, then we draw $P_{1}$ above the edge $c_{i-1}c_{i}$ and we draw the edge $e_{i}$ straight, without going around all drawings of $F_{1},\dots,F_{i-1}$. In this case $e_{i}$ is directed from $P_{1}$ to $P_{2}$. Hence, $e_{i}$ crosses the edge $u_{i-1}c_{i}$ if $u_{i-1}c_{i}$ is directed from $c_{i}$ to $u_{i-1}$; see the situation for $c_{2}$ in Fig. 7. Note that by Property 4 of Lemma 1, $e_{i}$ may receive at most a second crossing when we draw the remainder of $F_{i}$ in the usual way.
@@
Otherwise, that is, if $P_{1}$ contains an edge directed towards the first endpoint $w_{i}$ of $P_{1}$, let $f$ be the first such edge. We then place the part of $P_{1}$ up to the first endpoint of $f$ below the edge $c_{i-1}c_{i}$; see $w_{5}$ and $f$ in Fig. 7. If the edge $u_{i-1}c_{i}$ is directed towards $c_{i}$, we draw it between $w_{i}$ and the edge $c_{i-1}c_{i}$. Then it crosses the edge $c_{i-1}w_{i}$ but no other edge. We place the second endpoint of $f$ below all vertices in $V(F_{1})\cup\dots\cup V(F_{i-1})$ and continue with the remainder of $F_{i}$ as usual.
@@
In any case, if $1<i<k$, then the last vertex $u_{i-1}$ of $F_{i-1}$ is connected to $c_{i}$ and $c_{i-1}$ is connected to the first vertex $w_{i}$ in $F_{i}$. These two edges may cross each other; see the crossings highlighted in orange in Fig. 7. If the edge $c_{i-1}w_{i}$ goes, say, up but the following outer edges go down until a vertex $v_{k}$ below $c_{i}$ is reached, then the edge $c_{i-1}w_{i}$ may be crossed a second time by the edge $v_{k-1}v_{k}$; see the crossing labeled $x$ on the edge $c_{3}w_{4}$ in Fig. 7. Observe that in this case the path $P_{1}$ is the directed path from $w_{i}$ to $v_{k-1}$. Thus, due to Property 4 of Lemma 1, edge $v_{k-1}v_{k}$ had been crossed at most once within its fan. Also $c_{i-1}w_{i}$ cannot have a third crossing. Thus, all edges are crossed at most twice. ∎
@@
One can argue that every maximal pathwidth-2 graph can be generated from a maximal outerpath by connecting some pairs of adjacent vertices using an arbitrary number of (new) paths of length 2. In spite of the simplicity of this operation, we cannot hope to generalize the above result to pathwidth-2 graphs; see the linear lower bound on the upward local crossing number for such graphs stated in Theorem 2.
@@ 5 Testing Upward 1-Planarity
We start with a definition. Let $G_{1}$ and $G_{2}$ be any two st–digraphs. Let $s_{i}$ be the source and $t_{i}$ be the sink of $G_{i}$, with $i=1,2$. Let $G$ be a digraph that contains both $G_{1}$ and $G_{2}$ as induced subgraphs. Let $\Gamma$ be a drawing of $G$ and let $\Gamma_{1,2}$ be the drawing obtained by restricting $\Gamma$ to the vertices and edges of $G_{1}\cup G_{2}$. We say that $G_{1}$ and $G_{2}$ fully cross in $\Gamma$ if in $\Gamma_{1,2}$ every $s_{1}t_{1}$-path (i.e., a path directed from $s_{1}$ to $t_{1}$) crosses every $s_{2}t_{2}$-path. See Figs. 8(a) and 8(b) for examples of st-digraph $G_{1}$ and $G_{2}$ that do not fully cross or fully cross in a drawing of $\Gamma_{1,2}$, respectively.
@@
We now define a few gadgets; all of them are planar st-graphs. For positive integers $b$ and $q$, let a $(b,q)$-parallel be the parallel composition of $b$ oriented paths each consisting of $q$ edges; see Fig. 9(a). For a positive integer $p$, let a $(p)$-gate be the parallel composition of an oriented edge and a $(p-1,2)$-parallel; see Fig. 9(b). For positive integers $h$, $q$, and $a$, let an $(h,q,a)$-chain consist of a series of $h$ $(q)$-gates, followed by exactly one $(a)$-gate, followed again by $h$ $(q)$-gates; see Fig. 9(c).
@@
Conversely, assume that the instance $I$ of 3-Partition admits a solution. In order to prove that $G$ admits an upward 1-planar drawing, by hypothesis (ii) it suffices to construct an upward 1-planar drawing $\Gamma_{AB}$ of $G_{A}$ and $G_{B}$ where $G_{A}$ and $G_{B}$ fully cross. We enumerate the paths of $G_{B}$ as $\pi_{1},\pi_{2},\dots,\pi_{b}$. We also enumerate the $2(b-1)$ many $(W+1)$-gates and the $(a_{i})$-gate of a $(b-1,W+1,a_{i})$-chain as $g_{i,1},g_{i,2},\cdots,g_{i,2(b-1)}$, where $g_{i,1}$ contains the source $s_{A}$ of $G_{A}$ and $g_{i,2(b-1)}$ contains the sink $t_{A}$ of $G_{A}$ (see also Fig. 10). Consider a path $\pi_{j}$ with $1\leq j\leq b$, and let $\{a_{\chi},a_{\lambda},a_{\mu}\}$ be the $j$-th bin of the solution of $I$. Let $\nu$ be an index such that $1\leq\nu\leq b$.
@@
  • If $\nu$ is one of $\chi,\lambda,\mu$, then $\pi_{j}$ crosses the $(a_{\nu})$-gate of the $(b-1,W+1,a_{\nu})$-chain (see, for example, the red path crossing $g_{1,3}$ in Fig. 10).
  • If $\nu$ is not one of $\chi,\lambda,\mu$ and the path $\pi_{h}$ that crosses the $(a_{\nu})$-gate of the $(b-1,W+1,a_{\nu})$-chain is such that $h>j$, then $\pi_{j}$ crosses the gate $g_{\nu,j}$ of the $(W+1,b-1,a_{\nu})$-chain (see, for example, the blue path crossing $g_{2,2}$ in Fig. 10).
  • Otherwise ($h<j$), path $\pi_{j}$ crosses the gate $g_{\nu,b-1+j}$ of the $(b-1,W+1,a_{\nu})$-chain (see, for example, the red path crossing $g_{5,5}$ in Fig. 10).
By the procedure above, if $\pi_{j}$ crosses a gate $g(\nu,q)$, with $1\leq\nu\leq k$ and $1\leq q\leq 2(b-1)$, there is no other $\pi_{h}$, with $h\neq j$, such that $\pi_{h}$ crosses $g(\nu,q)$. Also, the number of edges that $\pi_{j}$ crosses is $a_{\chi}+a_{\lambda}+a_{\mu}+(k-3)(W+1)=W+(k-3)(W+1)$, which is the number of edges of $\pi_{j}$. Hence, it is possible to draw $\pi_{j}$ in such a way that each of its edges is crossed exactly once and no edge of $G_{A}$ is crossed more than once. This concludes the proof. ∎
Theorem 6.
@@
Proof.
It is immediate to observe that upward 1-planarity testing is in the NP class of complexity (one can guess an upward embedding and test it in polynomial time). We show that it is NP-hard. For each case in the statement it suffices to exhibit a digraph $G$ that contains $G_{A}$ and $G_{B}$ as sub-graphs and that satisfies the conditions of Theorem 5.